### v1

@@ 3.2. Design Goals
    Systematic In-the-Moment Reflexivity Support (DG1): To address Inconsistent Reflection Support (C1), which arises from inconsistent manual reflection and fragmented insight recording (B1, B2), we aimed to provide structured, in-the-moment reflexive prompts. This goal is to integrate reflection seamlessly into the coding workflow, anchoring insights to specific data without requiring researchers to switch contexts.
@@
    Automated Code Evolution Tracking (DG2): To address Untracked Code Evolution (C2), originated from labor-intensive documentation (B3) and reactive change management (B4), our objective was to automatically create a transparent, auditable history of code evolution. This includes proactive code drift detection to move beyond the reliance on manual researcher awareness (B4), making the entire analytical journey, including splits, merges, and definition edits, both visible and rigorous.
@@
    Collaborative Positionality Integration (DG3): To address Inaccessible Collaboration Perspectives (C3), a challenge deriving from reliance on informal self-disclosure (B5), separate discussions (B6), and sporadic prompting (B7), we aimed to make positionality an actionable and visible part of the analysis by surfacing interpretive differences and scaffolding dialogue around researchers’ unique lenses.
@@ 4. System Design — Reflexis
With the design goals in §3.2 (DG1–DG3) in mind, we implemented Reflexis, a web-based collaborative environment for reflexive thematic analysis that centers in-situ reflection, records code evolution for transparency, and makes positionality actionable. We first outline a usage scenario, then detail the core features, the prompt design that operationalizes our reflexive practices, and finally, implementation details.
@@ 4.1. Workflow & Usage Scenario
During independent coding, Alice encounters a city manager’s quote on feedback systems. Activating the ReflexiveLens, she uses the Positionality prompt to note that her government background reveals “a deep-seated institutional fear of failure,” and codes it Bureaucratic Hurdles. Bob, coding the same passage through his academic lens, notes that the system “privileges those with resources,” applying Participation Barrier.
@@ 4.2.1. ReflexiveLens: In-situ reflection and situated viewing
Figure 2. Reflexis: Supporting Reflexive Analysis. (1) User can easily engage in a just-in-time guided reflection during analysis, the reflection process is designed to be lightweight and optional, reducing labor and encouraging reflexivity. All users can review reflexive notes directly under the coding workspace, or (2) engage in additional reflection. (3) Users can also access reflexive notes within specific code on the sidebar and can filter by unique user, as well as the ability to create structured summary.
To support DG1, Reflexis promotes and surfaces reflexivity during key analytic moments. When a coder applies a code to a text snippet, or reflects on any prior analysis, the coder can engage the ReflexiveLens panel to add reflexive memos. Drawing from our formative study’s findings on inconsistent reflection support (see Section 3.1.1) and established reflexive practices (i.e., Probst and Berenson (2014)), the panel provides a series of pre-populated prompts to help anchor the coder’s analysis. These include prompts for:
@@
Complementing these structured prompts, a free-form notes section is also included to capture any emergent insights or unstructured reflections that fall outside the provided categories. The interaction is deliberately lightweight and optional, so that reflection occurs “just-in-time,” without interrupting the coding process. These reflexive memos are stored to maintain provenance between the excerpt and the applied code, ensuring that researchers can later reconstruct context.
@@ 4.2.2. Making Code Evolution Transparent and Tangible
Figure 3. Reflexis: Making Code Evolution Transparent and Tangible. To make the analytical process visible and consistent, Reflexis first (1) actively engages users with codes as evolving concepts by monitoring for code drift. When a shift in code application is detected, it alerts researchers, prompting a deliberate re-evaluation of the ongoing analysis and encouraging reflective practices. Supporting this reflective process, the system provides a detailed record of the analysis: the (2) code-level analysis history records every analytic event, including code creation, application, merges, and splits. Finally, (3) the project-level analysis history visualizes analytical provenance, tracing how codes are transformed over time.
To support DG2, Reflexis treats the codebook as a living object and records its evolution at two levels. At the code level, the system maintains a versioned code-level analysis history including creation, renames, definition edits, merges, and splits, with each user action timestamped and attributed to the researcher who executed it. This enables readers to understand not only what changed but also how the code came to be. At the project level, a global project-level analysis history aggregates events across codes and documents to support coordination (i.e., “what changed since our last meeting?”) and reporting (i.e., “how did the codebook evolve since the beginning of the project?”). Coders zoom in on the timeline to reveal details-on-demand, and they can export snapshots for auditing or producing methods and appendices in their final reports.
@@ 4.3. Prompt Design and LLM Usage
We adopt a narrow, specific role for language models in Reflexis: assist reflection and discussion without generating codes or suggesting what to apply. All prompt wordings and safety constraints are provided in Appendix A.2, and here we’ll only describe their function and scope.
Models are used in four features. A researcher can generate a digest of linguistic patterns, positionality effects, and articulated alternatives with reflexive stream structured summary. This allows researchers to gain a high-level overview of their team’s analytical approach and helps to surface emergent insights for further development. For positionality-aware discussion prompts, short discussion prompts are synthesized from the coded context and coders’ positionality profiles to help collaborators explain why interpretations may differ. The system does not adjudicate or prescribe one researcher’s choices over another. For code drift alert, the system compares the user-selected candidate corpus with recent exemplars from the same code, and flags potential broadening or misuse for human review. Finally, positionality keywords provides concise, LLM-simplified cues from a coder’s profile and recent memos, shown beside their highlights to keep backgrounds salient during discussion. Reflexis uses OpenAI’s GPT-5 and GPT-5-mini with low to medium reasoning effort.
@@ 4.4. System Implementation
The frontend is built with Next.js and Tailwind CSS, with the main coding workspace supporting character-precise highlights and overlapping annotations. Global analysis history is visualized with the react-flow component to show the project-level sequence of coding activity.
We use Google’s Firebase services for real-time collaboration: a Firestore database for real-time synchronization, anonymous authentication for easy onboarding and user identification, and managed hosting for backend deployment. The UI subscribes to shared documents, annotations, and profiles, so updates appear instantly across clients without manual refresh. To maintain transparency, every analytic action (apply, edit, rename, merge, split) is written to an immutable event log that provides history views and exports.
@@ 4.4.1. Percentage Agreement
Discussion Focus is calculated based on percentage agreement, a common metric in qualitative analysis (Miles and Huberman, 1994), using the following formula:
@@
We use this metric because common alternatives like Cohen’s Kappa can be statistically unsound when applied to more than two coders due to the violation of statistical assumptions (McHugh, 2012). We emphasize that this metric (and its color cue) is used solely to route attention to potentially fruitful discussion points, not to optimize for high agreement, consistent with our stance that principled divergence is often desirable in RTA, and we note that visual encodings can otherwise nudge behavior toward consensus (Leshed et al., 2010).
@@ 5.1. Participants, Setting, and Ethics
The majority of participants (10/12) are graduate students (i.e., master’s, PhD), one participant is a university professor, and another is working as a data analyst. Most of the participants come from an HCI background (10/12), with intersections spanning across NLP, visualization, and computing education. The rest of the participants came from a more technical background (i.e., software engineer, 2/12). On average, participants have 3.75 years of qualitative data analysis experience (Min: 1, Max: 7, Median: 4). All participants described Reflexive Thematic Analysis (RTA) as one of the methods they used to analyze qualitative data in the past. Participants reported mainly using spreadsheet software (i.e., Microsoft Excel, Google Sheet; 10/12) and text editors (i.e., Microsoft Word, Google Doc; 8/12) as part of their past RTA projects, with a handful of participants choosing QDA software like NVivo (5/12), ATLAS.ti (2/12), and Dedoose (1/12). Participants do not have a consensus for norms on reporting about their analytical practices or artifacts for the broader scientific community. Half of the participants indicated they do not share any artifacts, while three indicated they would share an end-of-study statement. Participants were asked to share their theoretical background and view on the topic of analysis (to be discussed below) without seeing the actual corpus, and were paired up to prioritize diverse perspectives, leading to the formation of 6 pairs (see Appendix A.4).
All sessions took place remotely. Each session lasted approximately 1.5-2 hours, and participants received a rate of \$20/hour compensation upon successful completion. The protocol, consent, and compensation were approved by our university IRB.
@@ 5.2. Study Design
In this study, participants were instructed to use Reflexis to perform reflexive thematic analysis (RTA) as a pair. Our research goal was to understand how Reflexis might reshape the deep, interpretive practices of collaborative qualitative analysis, as opposed to measure efficiency gains. Consequently, a direct comparison to a baseline tool would not effectively serve our research objective. We further explain our reasoning for not employing a standard baseline condition immediately after, in Section 5.2.1.
Our study protocol on collaborative reflexive analysis follows common qualitative guidance emphasizing (i) independent analytic passes, (ii) transparency and evolution of code definitions, and (iii) negotiated understanding via discussion. In particular, we refer to the process documented by O’Connor and Joffe (2020) (see Figure 3 in the corresponding paper). We randomly assigned one participant as Coder 1 and the other as Coder 2. Coder 1 first provides the initial code definitions and completes an initial pass on the provided corpus. Coder 2 then codes the same snippets using Coder 1’s generated code list, but without seeing which codes Coder 1 has applied to text segments. The pair subsequently reviews differences and rationales, mediated by system features that surface code overlaps, definition drift, and unmatched assignments.
Following prior evaluations of AI-supported qualitative coding tools that adopt intensive, single-session designs to study collaborative sense-making in depth (e.g., (Gebreegziabher et al., 2023; Gao et al., 2024)), our study similarly invites pairs to complete a bounded, collaborative RTA episode with Reflexis. While a 1.5-hour session cannot fully capture the multi-week qualitative analysis practices, it enables us to closely observe how analysts interact with and reason through Reflexis’s reflexivity-oriented features within a realistic, collaborative coding session.
@@ 5.2.1. Reasoning for Not Employing a Baseline Condition
Instead, our study was designed to observe how participants integrated Reflexis into their workflows, allowing us to gather rich, situated insights into its impact on their analytical process. This approach enables us to evaluate our system on its own terms, by its ability to support the core principles of reflexive analysis (Braun and Clarke, 2021).
@@ 5.2.2. Analysis Dataset
The analysis corpus used during the evaluation study is a synthetic, context-faithful set of interview transcripts derived from a prior CHI paper on “Contestable Camera Cars” (Alfrink et al., 2023). We chose this research topic because the original study employed reflexive thematic analysis, aligning with our methodology. In addition, the subject of “Contestable Camera Cars” is readily understandable and requires minimal domain-specific knowledge, making it highly accessible for interpretation. We prompted a large language model [^4] Given identical system prompts, we compared gemini-2.5-pro (thinking budget: 32k tokens) and gpt-5 (reasoning effort: high). We collectively elected to use the gemini-2.5-pro outputs, which yielded more realistic, conversational transcripts for our study. with the reference paper to generate three distinct interview transcripts that (i) stayed within the same civic-technology context, tone, and terminology; and (ii) echo the themes reported by the original authors. Participants were not informed that the transcripts were synthetic, and were instructed to analyze one of the three synthetic transcripts as a pair, in random order. This choice preserves ecological validity by retaining content and findings from a real study while avoiding privacy concerns and domain prerequisites. The full system prompt can be found in Appendix A.3.
@@ 5.3. Procedure
The session began with a 15-minute orientation. The moderator framed the upcoming activity as a typical Reflexive Thematic Analysis (RTA) pass and demonstrated the core features of Reflexis. Participants were briefed on the study’s specific collaborative workflow, which required sequential roles (Coder 1 followed by Coder 2) prior to a joint discussion, following process described by O’Connor and Joffe (2020).
Following orientation, participants engaged in an independent coding phase lasting approximately 30 minutes. Coder 1 worked first (approx. 15 minutes) to code assigned snippets, creating or renaming codes as necessary, writing short reflexive notes, and refining code definitions. Coder 2 then coded the same snippets (approx. 10 minutes) utilizing the code list and definitions developed by Coder 1. Furthermore, while Coder 2 utilized Coder 1’s schema, they were blinded to Coder 1’s specific text segment assignments to ensure independent application.
@@ 5.4. Measures and Analysis
We focused on qualitative feedback and perceived usefulness gathered through (i) brief pre/post-study surveys (open-ended items on reflexivity, fit with existing workflow, and perceived support), (ii) a semi-structured post-study interview, and (iii) in-study interactions (think-aloud comments, on-canvas/reflection notes, and mediated discussion logs). We also capture descriptive traces of the analysis process via screen recording, e.g., codebook edits (merges/splits/renames), definition revisions, and disagreement flags, to contextualize participants’ narratives.
We do not measure task time, “accuracy” of any kind (i.e., inter-rater reliability, participants’ accuracy against ‘ground truth’), or other aggregate quantitative metrics. In reflexive thematic analysis, higher agreement or faster completion does not necessarily indicate better analytic quality; such metrics can inadvertently suppress productive interpretive differences and reflexive engagement (McDonald et al., 2019). Accordingly, our traces are used to illuminate how interpretations evolve and are negotiated, not to compute efficiency or reliability scores.

### v2

@@ 3.2. Design Goals
    Systematic In-the-Moment Reflexivity Support (DG1): To address Inconsistent Reflection Support (C1), which arose from inconsistent manual reflection and fragmented insight recording (B1, B2), we aimed to provide structured, in-the-moment reflexive prompts. This goal aimed to integrate reflection seamlessly into the coding workflow, anchoring insights to specific data without requiring researchers to switch contexts.
@@
    Automated Code Evolution Tracking (DG2): To address Untracked Code Evolution (C2), originating from labor-intensive documentation (B3) and reactive change management (B4), our objective was to automatically create a transparent, auditable history of code evolution. This includes proactive Code Drift Alert to move beyond the reliance on manual researcher awareness (B4), making the entire analytical journey, including splits, merges, and definition edits, both visible and rigorous.
@@
    Collaborative Positionality Integration (DG3): To address Inaccessible Collaboration Perspectives (C3), a challenge stemming from reliance on informal self-disclosure (B5), separate discussions (B6), and sporadic prompting (B7), we aimed to make positionality an actionable and visible part of the analysis by surfacing interpretive differences and scaffolding dialogue around researchers’ unique lenses.
@@ 4. System Design — Reflexis
With the design goals in §3.2 (DG1–DG3) in mind, we implemented Reflexis, a web-based collaborative environment for reflexive thematic analysis that centers in-situ reflection, records code evolution for transparency, and makes positionality actionable. We first outline a usage scenario, then detail the core features, the prompt design that operationalizes our reflexive practices, and finally, technical details. We open-sourced the code for Reflexis at: https://github.com/harryye930/reflexis.
@@ 4.1. Workflow & Usage Scenario
During independent coding, Alice encounters a city manager’s quotation on feedback systems. Activating the ReflexiveLens, she uses the Positionality prompt to note that her government background reveals “a deep-seated institutional fear of failure,” and codes it Bureaucratic Hurdles. Bob, coding the same passage through his academic lens, notes that the system “privileges those with resources,” applying Participation Barrier.
@@ 4.2.1. ReflexiveLens: In-situ reflection and situated viewing
Figure 2. Reflexis: Supporting Reflexive Analysis. (1) Users can easily engage in a just-in-time guided reflection during analysis. The reflection process is designed to be lightweight and optional, reducing labor while encouraging reflexivity. All users can review reflexive notes directly under the coding workspace, or (2) engage in additional reflection. (3) Users can also access reflexive notes for a specific code within the sidebar and filter by unique user, in addition to creating structured summaries.
To support DG1, Reflexis promotes and surfaces reflexivity during key analytic moments. When a coder applies a code to a text snippet, or reflects on any prior analysis, the coder can use the ReflexiveLens panel to add reflexive memos. Drawing from our formative study’s findings on inconsistent reflection support (see Section 3.1.1) and established reflexive practices (e.g., Probst and Berenson (2014)), the panel provides a series of pre-populated prompts to help anchor the coder’s analysis. These include prompts for:
@@
Complementing these structured prompts, a free-form notes section is included to capture any emergent insights or unstructured reflections that fall outside the provided categories. The interaction is deliberately lightweight and optional, so that reflection occurs “just-in-time,” without interrupting the coding process.
@@ 4.2.2. Making Code Evolution Transparent and Tangible
Figure 3. Reflexis: Making Code Evolution Transparent and Tangible. To make the analytical process visible and consistent, Reflexis first (1) actively engages users with codes as evolving concepts by monitoring for code drift. When a shift in code application is detected, it alerts researchers, prompting a deliberate re-evaluation of the ongoing analysis and encouraging reflective practices. To support this reflective process, the system provides a detailed record of the analysis: the (2) code-level analysis history records every analytic event, including code creation, application, merges, and splits. Finally, (3) the project-level analysis history visualizes analytical provenance, tracing how codes are transformed over time.
To support DG2, Reflexis treats the codebook as a living object and records its evolution at two levels. At the code level, the system maintains a versioned code-level analysis history including creation, renames, definition edits, merges, and splits, with each user action timestamped and attributed to the user who executed it. This enables readers to understand not only what changed but also how the code came to be. At the project level, a global project-level analysis history aggregates events across codes and documents to support coordination (i.e., “what changed since our last meeting?”) and reporting (i.e., “how did the codebook evolve since the beginning of the project?”). Coders zoom in on the timeline to reveal details-on-demand, and they can export snapshots for auditing or producing methods and appendices in their final reports.
@@ 4.3. Prompt Design and LLM Usage
We adopted a narrow, specific role for language models in Reflexis: assist reflection and discussion without generating code or suggesting what to apply. All prompt wordings and safety constraints are provided in Appendix A.2; only function and scope are described here.
Models are used in four features. A researcher can generate a digest of linguistic patterns, positionality effects, and articulated alternatives with reflexive stream structured summary. This allows researchers to gain a high-level overview of their team’s analytical approach and helps to surface emergent insights for further development. For Positionality-Aware Discussion Prompts, short discussion prompts are synthesized from the coded context and coders’ positionality profiles to help collaborators explain why interpretations may differ. The system does not adjudicate or prescribe one researcher’s choices over another. For Code Drift Alert, the system compares the user-selected candidate corpus with recent exemplars from the same code, and flags potential broadening or misuse for human review. Finally, Positionality Keywords provide concise, LLM-simplified cues from a coder’s profile and recent memos, shown beside their highlights to keep backgrounds salient during discussion. Reflexis uses OpenAI’s GPT-5 and GPT-5-mini with low to medium reasoning effort.
@@ 4.4. System Implementation
The frontend was built with Next.js and Tailwind CSS, with the main coding workspace supporting character-precise highlighting and overlapping annotations. Project-level analysis history was visualized using the React Flow library.
We used Google’s Firebase services for real-time collaboration. In particular, the Firestore database was used for real-time synchronization, anonymous authentication for easy onboarding and user identification, and managed hosting for backend deployment. The UI subscribed to shared documents, annotations, and profiles, so updates appeared instantly across clients without manual refresh. To maintain transparency, every analytic action (apply, edit, rename, merge, split) was written to an immutable event log that provides history views and exports.
@@ 4.4.1. Percentage Agreement
The agreement metric for Discussion Focus is calculated based on percentage agreement, a common metric in qualitative analysis (Miles and Huberman, 1994), using the following formula:
@@
We used this metric because common alternatives like Cohen’s Kappa can be statistically unsound when applied to more than two coders due to the violation of statistical assumptions (McHugh, 2012). We emphasize that this metric (and its color cue) is used solely to route attention to potentially fruitful discussion points, not to optimize for high agreement, consistent with our stance that principled divergence is often desirable in RTA. Yet, we also note the possibility that visual encodings may nudge behavior toward consensus (Leshed et al., 2010).
@@ 5.1. Participants, Setting, and Ethics
The majority of participants (10/12) were graduate students (i.e., master’s, PhD), one participant was a university professor, and another worked as a data analyst. Most of the participants came from an HCI background (10/12), with intersections spanning across NLP, visualization, and computing education. The rest of the participants had a more technical background (i.e., software engineers, 2/12). On average, participants had 3.75 years of qualitative data analysis experience (Min: 1, Max: 7, Median: 4). All participants described Reflexive Thematic Analysis (RTA) as one of the methods they used to analyze qualitative data in the past. Participants reported mainly using spreadsheet software (i.e., Microsoft Excel, Google Sheets; 10/12) and text editors (i.e., Microsoft Word, Google Docs; 8/12) as part of their past RTA projects, with a handful of participants choosing QDA software like NVivo (5/12), ATLAS.ti (2/12), and Dedoose (1/12). Participants did not have a consensus for norms on reporting about their analytical practices or artifacts for the broader scientific community. Half of the participants indicated they did not share any artifacts, while three indicated they would share an end-of-study statement. Participants were asked to share their theoretical background and view on the topic of analysis (discussed below) without seeing the actual corpus, and were paired up to prioritize diverse perspectives, leading to the formation of 6 pairs (see Appendix A.4).
All sessions took place remotely. Each session lasted approximately 1.5-2 hours, and participants received compensation at a rate of \$20/hour upon successful completion. The protocol, consent, and compensation were approved by our university IRB office (protocol # 49033).
@@ 5.2. Study Design
In this study, participants were instructed to use Reflexis to perform reflexive thematic analysis (RTA) as a pair. Our research goal was to understand how Reflexis might reshape the deep, interpretive practices of collaborative qualitative analysis, as opposed to measuring efficiency gains. Consequently, a direct comparison to a baseline tool would not effectively serve our research objective. We further explain our reasoning for not employing a standard baseline condition in Section 5.2.1.
Our study protocol on collaborative reflexive analysis followed common qualitative guidance emphasizing (i) independent analytic passes, (ii) transparency and evolution of code definitions, and (iii) negotiated understanding via discussion. In particular, we referred to the process documented by O’Connor and Joffe (2020) (see Figure 3 in the corresponding paper). We randomly assigned one participant as Coder 1 and the other as Coder 2. Coder 1 first provided the initial code definitions and completed an initial pass on the provided corpus. Coder 2 then coded the same snippets using Coder 1’s generated code list, but without seeing which codes Coder 1 had applied to text segments. The pair subsequently reviewed differences and rationales, mediated by system features that surface code overlaps, definition drift, and unmatched assignments.
Following prior evaluations of AI-supported qualitative coding tools that adopt intensive, single-session designs to study collaborative sense-making in depth (e.g., (Gebreegziabher et al., 2023; Gao et al., 2024)), our study similarly invites pairs to complete a bounded, collaborative RTA episode with Reflexis. While a 1.5-hour session cannot fully capture a multi-week qualitative analysis practice, it enables us to closely observe how analysts interact with and reason through Reflexis’s reflexivity-oriented features within a realistic, collaborative coding session.
@@ 5.2.1. Reasoning for Not Employing a Baseline Condition
Instead, our study was designed to observe how participants integrated Reflexis into their workflows, allowing us to gather rich, situated insights into its impact on their analytical process. This approach enables us to evaluate our system on its own terms via its ability to support the core principles of reflexive analysis (Braun and Clarke, 2021).
@@ 5.2.2. Analysis Dataset
The analysis corpus used during the evaluation study was a synthetic, context-faithful set of interview transcripts derived from a prior CHI paper on “Contestable Camera Cars” (Alfrink et al., 2023). We chose this research topic because the original study employed reflexive thematic analysis, aligning with our methodology. In addition, the subject of “Contestable Camera Cars” is readily understandable and requires minimal domain-specific knowledge, making it highly accessible for interpretation. We prompted a large language model [^4] Given identical system prompts, we compared gemini-2.5-pro (thinking budget: 32k tokens) and gpt-5 (reasoning effort: high). We collectively elected to use the gemini-2.5-pro outputs, which yielded more realistic, conversational transcripts for our study. with the reference paper to generate three distinct interview transcripts that (i) preserved the tone and terminology while staying within the civic-technology context; and (ii) echo the themes reported by the original authors. Participants were not informed that the transcripts were synthetic, and were instructed to analyze one of the three synthetic transcripts as a pair, in random order. This choice preserved ecological validity by retaining content and findings from a real study while avoiding privacy concerns and domain prerequisites. The full system prompt is available in Appendix A.3.
@@ 5.3. Procedure
The session began with a 15-minute orientation. The moderator framed the upcoming activity as a typical RTA pass and demonstrated the core features of Reflexis. Participants were briefed on the study’s specific collaborative workflow, which required sequential roles (Coder 1 followed by Coder 2) prior to a joint discussion, following the process described by O’Connor and Joffe (2020).
Following orientation, participants engaged in an independent coding phase lasting approximately 30 minutes. Coder 1 worked first (approx. 15 minutes) to code assigned snippets, creating or renaming codes as necessary, writing short reflexive notes, and refining code definitions. Coder 2 then coded the same snippets (approx. 10 minutes) utilizing the code list and definitions developed by Coder 1. Furthermore, while Coder 2 utilized Coder 1’s schema, they were blinded to Coder 1’s specific text segment code assignments to ensure independent application.
@@ 5.4. Measures and Analysis
We focused on qualitative feedback and perceived usefulness gathered through (i) brief pre/post-study surveys (open-ended items on reflexivity, fit with existing workflow, and perceived support), (ii) a semi-structured post-study interview, and (iii) in-study interactions (think-aloud comments, on-canvas/reflection notes, and mediated discussion logs). We also captured descriptive traces of the analysis process via screen recording, e.g., codebook edits (merges/splits/renames), definition revisions, and disagreement flags, to contextualize participants’ narratives.
We did not measure task time, “accuracy” of any kind (i.e., inter-rater reliability, participants’ accuracy against ‘ground truth’), or other aggregate quantitative metrics. In RTA, higher agreement or faster completion does not necessarily indicate better analytic quality; as such metrics can inadvertently suppress productive interpretive differences and reflexive engagement (McDonald et al., 2019). Accordingly, our measurements were selected to illuminate how interpretations evolve and are negotiated, not to compute efficiency or reliability scores.